OpenStudy (anonymous):

Simplify: –4(x – 5) + 3(a – 7) A. –4x + 3a – 1 B. –4x + 3a – 41 C. –4x + 3a + 1 D. –4x + 3a + 41

OpenStudy (luigi0210):

Distribute

OpenStudy (anonymous):

.. o.0

OpenStudy (anonymous):

–4(x – 5) + 3(a – 7) You have to "distribute" the -4 into the (x+5) and the 3 into the (a-7) The distribute property states that a(x+y)=ax+by. For example, 9(1+2)=(9)(1)+(9)(2) Or, 6(5+x)=(6)(5)+(6)(x).

OpenStudy (anonymous):

um..I still don't get it but ok '.'

OpenStudy (jadeishere):

–4(x – 5) + 3(a – 7) they mean make it look like this (-4 * x) + (-4 * -20) + (3 * a) + (3*-7)

OpenStudy (anonymous):

Do I add that?

OpenStudy (jadeishere):

I don't know how to explain it well multiply everything inside of the parentheses example 2(b + 6) 2b + 12 see what I did there?

OpenStudy (anonymous):

ya

OpenStudy (jadeishere):

Do the same thing with –4(x – 5) + 3(a – 7) -4x + 20 then do the same with the other one [3(a - 7)]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would it be 21a - 7..?

OpenStudy (jadeishere):

so [-4(x -5)] + [3(a-7)] and you'll get the answer. If you still don't get it i'll walk you through it some more

OpenStudy (anonymous):

Would it be B..?

OpenStudy (jadeishere):

-4x + 3a – 41 let's see [-4(x -5)] + [3(a-7)] -4x + 20 + 3a - 21 add like terms, you tell me, is it?

OpenStudy (jadeishere):

if you look closely at 20 and - 21, it would not equal -41. So.. No, it is not B

OpenStudy (anonymous):

oh

OpenStudy (jadeishere):

you got the (-)(+) signs mixed up

OpenStudy (jadeishere):

when you multiply a - by a - it turns into a positive

OpenStudy (jadeishere):

in this case, you got -20 - 21 right? try it as 20 - 21

OpenStudy (anonymous):

1 so it's either A or C right?

OpenStudy (jadeishere):

is 20 - 21 a positive?

OpenStudy (anonymous):

yes..I believe it is, is the 21 a -21 or are they both positive?

OpenStudy (jadeishere):

it is a -21. 20 - 21 is a negative. since there is more of 21 than there is 20 it is -1

OpenStudy (anonymous):

ohh ok, I was confused at first whether it was a positive or negative but now I get it :)

OpenStudy (luigi0210):

Nice job Jadey~

OpenStudy (the_fizicx99):

Hm, \(\ \sf \Large a(b + c) \rightarrow a(b) + a(b) = ab + ac \) \(\ \sf \Large \color{blue}{–4}(\color{orange}{x} –\color{yellowgreen}{5}) \longrightarrow \color{blue}{–4}(\color{orange}{x}) = \color{red}{(-4x)}\) and \(\ \sf \Large \color{blue}{–4}(–\color{yellowgreen}{5}) = \color{brown}{(20)}\) \(\ \sf \Large \color{purple}{+3}(\color{aqua}{a} –\color{pink}{7}) \longrightarrow \color{purple}{3}(\color{aqua}{a}) = (3a)\) ~ and ~ \(\ \sf \Large \color{purple}{3}(–\color{pink}{7}) = (-21) \) \(\ \sf \Large \color{red}{(-4x)} + \color{brown}{(20)} + 3a + (-21) = 3a -4x -1\)

OpenStudy (the_fizicx99):

X_X Typo ;_;